OpenStudy (anonymous):

11.The formula for the perimeter P of a rectangle with length l and width w is P = 2l + 2w. Which of the following is a formula for the length of a rectangle in terms of the perimeter and width?

OpenStudy (anonymous):

A. L= P-w/2 B. L= P+w/2 C. L= P-2w/2 D. L= P+2W/2

OpenStudy (anonymous):